OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (4, –6) and is parallel to y = –1/2x + 5. (4 points) y = –2x + 2 y = 2x – 14 y = 1/2x – 8 y = –1/2x – 4

OpenStudy (ajprincess):

y=mx +c m-slope c-y_intercept For parallel lines the slopes are same. Does that hep @alexashea1997?

OpenStudy (anonymous):

kinda.

OpenStudy (ajprincess):

If u feel that u need a bit more help u can ask me.:) you knw the slope and a point is given. using both of them u can find c. y=mx+c -6=-(1/2)*(4)+c Find c. Then plug in the values of m and c in y=mx+c to find the equation

OpenStudy (anonymous):

thats better thanks!

OpenStudy (anonymous):

can you help me with more ?

OpenStudy (ajprincess):

ya sure

OpenStudy (anonymous):

17. Which equation below is not an equation that represents the graph shown below? https://www.connexus.com/content/media/456845-752011-21644-PM-1719707532.png y – 3 = 2(x – 6) y + 3 = (x + 6) 2y = x x – 2y = 0

OpenStudy (anonymous):

i think its either a or b

OpenStudy (ajprincess):

I am sorry. I am unable to view ur attachment. It asks for login

OpenStudy (anonymous):

can you see that ?